OpenStudy (anonymous):

What is the volume of a cube with edge 2x ft.

Directrix (directrix):

The volume of a cube formula is similar to that of the rectangular solid. For length, width, and height, the measure will be the same: (2x).

OpenStudy (anonymous):

I got 4x squared is that right?

OpenStudy (callisto):

Nope. Volume of cube = (length)^3

Directrix (directrix):

@DustinMHayes V = (length)^3 = (2x)^3 = (2x)*(2x)*(2x) V = ?